# Arbitrage-free catastrophe reinsurance valuation for compound dynamic contagion claims

### v1

@@ Abstract
In this paper, we consider catastrophe stop-loss reinsurance valuation for a reinsurance company with dynamic contagion claims. To deal with conventional and emerging catastrophic events, we propose the use of a compound dynamic contagion process for the catastrophic component of the liability. Under the premise that there is an absence of arbitrage opportunity in the market, we obtain arbitrage-free premiums for these contacts. To this end, the Esscher transform is adopted to specify an equivalent martingale probability measure. We show that reinsurers have various ways of levying the security loading on the net premiums to quantify the catastrophic liability in light of the growing challenges posed by emerging risks arising from climate change, cyberattacks, and pandemics. We numerically compare arbitrage-free catastrophe stop-loss reinsurance premiums via the Monte Carlo simulation method. Sensitivity analyzes are performed by changing the Esscher parameters and the retention level.
@@ 1 Introduction
The estimated cost of claims related to February and March 2022 South-East Queensland and New South Wales floods, which is the third most costly extreme weather event ever recorded in Australia, by the Insurance Council of Australia (ICA) is AU\$4.8 billion. Insurers’ losses from October 2022 Hurricane Ian, one of the most powerful storms ever recorded in the US, are expected to be between US\$42 billion and US\$57 billion. In July 2023, we have seen extreme weather events, i.e., floods in Asia, wildfires in Canada and deadly heatwaves in southern Europe and US, which can impact economies badly with huge financial cost.
Hurricane Helene slammed into Florida’s Gulf cost caused deadly flooding, dumping almost unprecedented amounts of rain in September 2024. AccuWeather Inc. estimated that total damage and economic loss could be between US\$145 billion and US\$160 billion. Spain has been affected by record-breaking rainfall and severe flooding in Valencia in late October 2024. The Southern California wildfires of January 2025 illustrate the dynamics of claims. Moody’s Risk Management Solutions (RMS) event response estimates that insured losses for Los Angeles firestorm events will likely range between US\$20 billion and US\$30 billion. These events present significant challenges to the financial stability of insurers and reinsurers, highlighting the need for improved models to predict claims from catastrophic events.
In May 2021 a ransomware cyberattack halted the US Colonial Pipeline by disrupting its computerized control systems. Nearly half of the US East Coast’s fuel supply depends on this pipeline, which transports gasoline, diesel, jet fuel and other refined products. The Colonial Pipeline Company acknowledged that it paid a ransom of \$US4.4 million worth of bitcoin, though the US Department of Justice was able to recover 64 bitcoins worth \$US2.3 million.  According to Travel, Logistics & Infrastructure article, airlines suffered \$168 billion in economic losses in 2020, although the COVID-19 pandemic was entering its endemic stages in some parts of the world (Bouwer et al., 2022).
Insurers provide coverage to individuals and organizations by underwriting policies that protect against economic and financial losses resulting from catastrophic events. However, insurers cannot cover all losses, as extreme risks pose a challenge to their financial viability. To ensure being solvent in case of exceptional huge losses, insurers transfer their risk portfolios by purchasing catastrophe reinsurance contracts.
@@
As a result, alternative point processes are needed to model claim and loss arrivals from both conventional and emerging catastrophic events, ensuring effective pricing of catastrophe insurance and reinsurance contracts. To articulate the ongoing challenge and complexity of new risk dynamics, we use a compound dynamic contagion process (CDCP) (Dassios and Zhao, 2011) for the catastrophic component of the liability of a reinsurer, (i.e. the aggregate catastrophe loss covered by reinsurance contract), denoted by $C_{t}$, where $N_{t}$ is the dynamic contagion process (DCP) which is a point process. This makes our study unique, and to our knowledge, using it for the catastrophic component of the liability of $C_{t}$ is the first contribution in the context of pricing for catastrophe reinsurance contract. For references on this compound point process, we refer you to Jang and Oh (2021) and relevant works therein.
To ensure risk-neutrality, we use an equivalent martingale probability measure $\mathbb{P}^{\ast}$ for the compound dynamic contagion process $C_{t}$, derived using the Esscher transform, to evaluate an arbitrage-free premium. In the field of actuarial science, risk-neutral pricing via the Esscher transform (Esscher, 1932) can be found, for example, in Gerber and Shiu (1994), Gerber and Shiu (1996), Bühlmann et al. (1996), Dassios and Jang (2003), Jang and Krvavych (2004), Siu et al. (2004), and Siu (2005).
This paper considers the fair valuation of catastrophe stop-loss reinsurance contracts. We arrange the paper as follows. In Section 2, we provide a mathematical definition for the compound DCP for $C_{t}$, which was introduced by Jang and Oh (2021). The infinitesimal generator of the joint process is also provided extending it to the time-inhomogeneous case, together with required moments. We describe our insurance market and discuss the condition of the absence of arbitrage in Section 3. Section 4 provides arbitrage-free catastrophe reinsurance premiums numerically obtained using the Monte Carlo simulation method. Sensitivity analyzes are performed by changing the Esscher parameters and the retention level. Section 5 concludes the paper. Appendix A details our Monte Carlo algorithms.
@@ 2.1 Dynamic contagion process
A definition for the standard dynamic contagion process (DCP) with time-homogeneous parameters will be provided by Definition 2.1, while a definition for the associated CDCP will be provided by Definition 2.2. We will extend to a version with time-inhomogeneous parameters later.
@@
Let $\{M_{t}\}_{t\in{\cal T}}$ be a Poisson point process on $(\Omega,{\cal F},\mathbb{P})$ such that its intensity process is given by $\rho>0$, for all $t\in{\cal T}$, $\mathbb{P}$-a.s.. Write $\{T_{1,i}\}_{i=1,2,\ldots}$ and $\{X_{i}\}_{i=1,2,\ldots}$ for the random jump times and sizes for the Poisson point process $\{M_{t}\}_{t\in{\cal T}}$, respectively. Again, suppose that $\{X_{i}\}_{i=1,2,\ldots}$ is a sequence of i.i.d. positive random variables on $(\Omega,{\cal F},\mathbb{P})$ with the common distribution $H(x)$, where $x>0$. Assume that under the probability measure $\mathbb{P}$, $\{X_{i}\}_{i=1,2,\ldots}$, $\{T_{1,i}\}_{i=1,2,\ldots}$ and $\{Y_{j}\}_{j=1,2,\ldots}$ are independent with each other.
@@
  • the initial intensity at time $t=0$, $\lambda_{0}>0$, $\mathbb{P}$-a.s.;
  • the constant mean-reverting level $a\geq 0$;
@@
${\cal F}^{\lambda}_{t}$ $:=\sigma\{\lambda_{u}|u\in[0,t]\}\vee{\cal N},$
${\cal F}^{N}_{t}$ $:=\sigma\{N_{u}|u\in[0,t]\}\vee{\cal N},$
where ${\cal N}$ is a collection of the $\mathbb{P}$-null subsets of ${\cal F}$; ${\cal F}_{1}\vee{\cal F}_{2}$ is the minimal $\sigma$-field containing the two $\sigma$-fields ${\cal F}_{1}$ and ${\cal F}_{2}$. Note that for all $t\in{\cal T}$, ${\cal F}^{N}_{t}\subset{\cal F}^{\lambda}_{t}$.
@@ 2.2 Compound dynamic contagion process
For the standard DCP with time-homogeneous parameters, the infinitesimal generator ${\cal A}$ of the joint process $\{\left(\lambda_{t},N_{t},C_{t},{M_{t}},{\Lambda_{t}},t\right)\}_{t\in{\cal T}}$ acting on a function $f\left(\lambda,n,c,{m},{\Lambda},t\right)\in\mathcal{D}\left(\mathcal{A}\right)$ is given by:
$\mathcal{A}f\left(\lambda,\Lambda,n,c,{m},{\Lambda},t\right)$
@@
$\quad+\lambda\left[\int_{0}^{\infty}\int_{0}^{\infty}f\left(\lambda+y,n+1,c+\xi,{m},{\Lambda},t\right)\mathop{}\!\mathrm{d}G(y)\mathop{}\!\mathrm{d}J(\xi)-f\left(\lambda,n,c,{m},{\Lambda},t\right)\right]$
$\quad+\rho\left[\int_{0}^{\infty}f\left(\lambda+x,n,c,{m+1},{\Lambda},t\right)\mathop{}\!\mathrm{d}H(x)-f\left(\lambda,n,c,{m},{\Lambda},t\right)\right],$ (2.2)
where $\mathcal{D}\left(\mathcal{A}\right)$ is the domain of the generator $\mathcal{A}$ such that $f\left(\lambda,n,c,{m},{\Lambda},t\right)$ is differentiable with respect to $\lambda$ and $t$, and
$\left|\int_{0}^{\infty}\int_{0}^{\infty}f\left(\lambda+y,n+1,c+\xi,{m},{\Lambda},t\right)\mathop{}\!\mathrm{d}G(y)\mathop{}\!\mathrm{d}J(\xi)-f\left(\lambda,n,c,{m},{\Lambda},t\right)\right|<\infty\text{,}$
$\left|\int_{0}^{\infty}f\left(\lambda+x,n,c,{m+1},{\Lambda},t\right)\mathop{}\!\mathrm{d}H(x)-f\left(\lambda,n,c,{m},{\Lambda},t\right)\right|<\infty\text{.}$
Based on Section 2.2, we can easily obtain the moments of $\lambda_{t}$, $N_{t}$ and $C_{t}$. To do so, we denote the first-order moments of $X$, $Y$ and $\Xi$, respectively, by
@@
$\mathbb{E}\left[\lambda_{t}\mid\lambda_{0}\right]~{}=~{}\left\{\begin{array}[]{ll}\frac{\mu_{H}\rho+a\delta}{\kappa}+\left(\lambda_{0}-\frac{\mu_{H}\rho+a\delta}{\kappa}\right)\mathrm{e}^{-\kappa t},&\kappa\neq 0,\\ \lambda_{0}+\left(\mu_{H}\rho+a\delta\right)t,&\kappa=0.\end{array}\right.$
where $\kappa:=\delta-\mu_{G}$. Under the stationary condition $\kappa>0$, the asymptotic first moment of $\lambda_{t}$ under $\mathbb{P}$ are given by
$\mu_{1}\ :=\ \lim_{t\rightarrow\infty}\mathbb{E}\left[\lambda_{t}\mid\lambda_{0}\right]\ =\ \frac{\mu_{H}\rho+a\delta}{\kappa}.$
@@
$\mathbb{E}\left[N_{t}\mid\lambda_{0}\right]~{}=~{}\left\{\begin{array}[]{ll}\mu_{1}t+\left(\lambda_{0}-\mu_{1}\right)\frac{1}{\kappa}\left(1-\mathrm{e}^{-\kappa t}\right),&\kappa\neq 0,\\ \lambda_{0}t+\frac{1}{2}\left(\mu_{H}\rho+a\delta\right)t^{2},&\kappa=0.\end{array}\right.$
@@
$\mathbb{E}\left[C_{t}\mid\lambda_{0}\right]~{}=~{}\left\{\begin{array}[]{ll}\mu_{J}[\mu_{1}t+\left(\lambda_{0}-\mu_{1}\right)\frac{1}{\kappa}\left(1-\mathrm{e}^{-\kappa t}\right)],&\kappa\neq 0,\\ \mu_{J}[\lambda_{0}t+\frac{1}{2}\left(\mu_{H}\rho+a\delta\right)t^{2}],&\kappa=0.\end{array}\right.$
2.3 Time-inhomogeneous compound dynamic contagion process
The parameters of the previously described processes are all time-homogeneous. More generally, we can extend to the time-inhomogeneous case by allowing the parameters to vary over time (except the rate of decay $\delta$). More precisely, the mean-reverting level $a(t)$ and the rate of externally-excited jump arrival $\rho(t)$ are deterministic functions of time $t$, and are bounded on all intervals $[0,t)$ (no explosions). The distribution function of the externally-excited jump sizes at any time $t$ is $H(x;t)$ for $x>0$ with $\mu_{H}(t)=\int_{0}^{\infty}x\,\mathop{}\!\mathrm{d}H(x;t)$. The distribution function of the self-excited jump sizes at any time $t$ is $G(y;t)$ for $y>0$ with $\mu_{G}(t)=\int_{0}^{\infty}y\,\mathop{}\!\mathrm{d}G(y;t)$. We assume that $a(t)$, $\rho(t)$, $G(y;t)$ and $H(y;t)$ are all Riemann integrable functions of $t$ and are all positive. The reason for this generalization will become apparent later in Section 3 where we carry out change of measure.

### v2

@@ Abstract
In this paper, we consider catastrophe stop-loss reinsurance valuation for a reinsurance company with dynamic contagion claims. To deal with conventional and emerging catastrophic events, we propose the use of a compound dynamic contagion process for the catastrophic component of the liability. Under the premise that there is an absence of arbitrage opportunity in the market, we obtain arbitrage-free premiums for these contracts. To this end, the Esscher transform is adopted to specify an equivalent martingale probability measure. We show that reinsurers have various ways of levying the security loading on the net premiums to quantify the catastrophic liability in light of the growing challenges posed by emerging risks arising from climate change, cyberattacks, and pandemics. We numerically compare arbitrage-free catastrophe stop-loss reinsurance premiums via the Monte Carlo simulation method. We also compare them with those from generalised compound Hawkes/compound Cox cases. Sensitivity analyses are performed by changing the retention level, the Esscher parameters and the intensity parameters.
@@ 1 Introduction
The estimated cost of claims related to February and March 2022 South-East Queensland and New South Wales floods, which is the third most costly extreme weather event ever recorded in Australia, by the Insurance Council of Australia (ICA) is AU\$4.8 billion. Insurers’ losses from October 2022 Hurricane Ian, one of the most powerful storms ever recorded in the US, are expected to be between US\$42 billion and US\$57 billion. In July 2023, we have seen extreme weather events, i.e., floods in Asia, wildfires in Canada, and deadly heatwaves in southern Europe and the US, which can impact economies badly with huge financial cost.
Hurricane Helene slammed into Florida’s Gulf coast causing deadly flooding, dumping almost unprecedented amounts of rain in September 2024. AccuWeather Inc. estimated that total damage and economic loss could be between US\$145 billion and US\$160 billion. Spain has been affected by record-breaking rainfall and severe flooding in Valencia in late October 2024. The Southern California wildfires of January 2025 illustrate the dynamics of claims. Moody’s Risk Management Solutions (RMS) event response estimates that insured losses for Los Angeles firestorm events will likely range between US\$20 billion and US\$30 billion. These events present significant challenges to the financial stability of insurers and reinsurers, highlighting the need for improved models to predict claims from catastrophic events.
In May 2021 a ransomware cyberattack halted the US Colonial Pipeline by disrupting its computerised control systems. Nearly half of the US East Coast’s fuel supply depends on this pipeline, which transports gasoline, diesel, jet fuel and other refined products. The Colonial Pipeline Company acknowledged that it paid a ransom of US\$4.4 million worth of bitcoin, though the US Department of Justice was able to recover 64 bitcoins worth US\$2.3 million. According to Travel, Logistics & Infrastructure article, airlines suffered \$168 billion in economic losses in 2020, although the COVID-19 pandemic was entering its endemic stages in some parts of the world (Bouwer et al., 2022).
Insurers provide coverage to individuals and organisations by underwriting policies that protect against economic and financial losses resulting from catastrophic events. However, insurers cannot cover all losses, as extreme risks pose a challenge to their financial viability. To ensure being solvent in case of exceptional huge losses, insurers transfer their risk portfolios by purchasing catastrophe reinsurance contracts.
@@
As a result, alternative point processes are needed to model claim and loss arrivals from both conventional and emerging catastrophic events, ensuring effective pricing of catastrophe insurance and reinsurance contracts. To articulate the ongoing challenge and complexity of new risk dynamics, we use a compound dynamic contagion process (CDCP) (Dassios and Zhao, 2011) for the catastrophic component of the liability of a reinsurer, (i.e., the aggregate catastrophe loss covered by reinsurance contract), denoted by $C_{t}$. This makes our study unique, and to our knowledge, using it for the catastrophic component of the liability of $C_{t}$ is the first contribution in the context of pricing for catastrophe reinsurance contract. For references on this compound point process, we refer to Jang and Oh (2021) and relevant citations therein. Recently, Liu and Jang (2025) applied enhanced dynamic contagion processes to study optimal asset allocation and reinsurance problems.
The CDCP is particularly well suited to modeling catastrophe reinsurance liabilities because its probabilistic structure directly reflects key empirical features of catastrophic loss arrivals. First, the self-exciting component captures claim clustering and loss amplification which are often observed after major catastrophes. Specifically, an initial event, (e.g., a severe flood or wildfire), generates subsequent related claims through infrastructure damage, supply-chain disruption, or aftershocks. Second, the external-excitation component allows for sudden exogenous shocks, such as hurricanes, pandemics, or cyber events, whose arrival times are not triggered by past insurance losses, but by external environmental or systemic factors. Third, the mean-reverting intensity ensures that periods of elevated risk eventually decay, consistent with the transient nature of catastrophe-driven claim activities.
To ensure risk-neutrality, we use an equivalent martingale probability measure $\tilde{\mathbb{P}}$ for the compound dynamic contagion process $C_{t}$, derived using the Esscher transform, to evaluate an arbitrage-free premium. In the field of actuarial science, risk-neutral pricing via the Esscher transform (Esscher, 1932) can be found, for example, in Gerber and Shiu (1994), Gerber and Shiu (1996), Bühlmann et al. (1996), Dassios and Jang (2003), Jang and Krvavych (2004), Siu et al. (2004), and Siu (2005).
This paper considers the fair valuation of catastrophe stop-loss reinsurance contracts. We arrange the paper as follows. In Section 2, we provide a mathematical definition for the CDCP for $C_{t}$, which was introduced by Jang and Oh (2021). The infinitesimal generator of the joint process is also provided extending it to the time-inhomogeneous case, together with required moments. We describe our insurance market and discuss the condition of the absence of arbitrage in Section 3. Section 4 provides arbitrage-free catastrophe reinsurance premiums numerically obtained using the Monte Carlo simulation method. Comparisons are also made with the generalised Hawkes case and the Cox case, respectively. Sensitivity analyses are performed by changing the retention level, the Esscher parameters and the intensity parameters. Section 5 concludes the paper. Appendix A defines an equivalent martingale measure under a non-zero interest rate, while Appendix B details our Monte Carlo algorithms.
@@ 2.1 Dynamic contagion process
A definition for the standard dynamic contagion process (DCP) with time-homogeneous parameters will be provided by Definition 2.1, while a definition for the associated CDCP will be provided by Definition 2.2. We will extend it to a version with time-inhomogeneous parameters later.
@@
Let $\{M_{t}\}_{t\in{\cal T}}$ be a Poisson point process on $(\Omega,{\cal F},\mathbb{P})$ such that its intensity process is given by $\rho>0$, for all $t\in{\cal T}$, $\mathbb{P}$-a.s. Write $\{T_{1,i}\}_{i=1,2,\ldots}$ and $\{X_{i}\}_{i=1,2,\ldots}$ for the random jump times and sizes for the Poisson point process $\{M_{t}\}_{t\in{\cal T}}$, respectively. Again, suppose that $\{X_{i}\}_{i=1,2,\ldots}$ is a sequence of i.i.d. positive random variables on $(\Omega,{\cal F},\mathbb{P})$ with the common distribution $H(x)$, where $x>0$. Assume that under the probability measure $\mathbb{P}$, $\{X_{i}\}_{i=1,2,\ldots}$, $\{T_{1,i}\}_{i=1,2,\ldots}$ and $\{Y_{j}\}_{j=1,2,\ldots}$ are independent with each other.
@@
  • the constant mean-reverting (baseline) intensity level $a\geq 0$;
  • the constant initial intensity at time $t=0$, $\lambda_{0}\geq a$, $\mathbb{P}$-a.s.;
@@
${\cal F}^{\lambda}_{t}$ $=\sigma\{\lambda_{u}\,|\,u\in[0,t]\}\vee{\cal N},$
${\cal F}^{N}_{t}$ $=\sigma\{N_{u}\,|\,u\in[0,t]\}\vee{\cal N},$
[…]
@@ 2.2 Compound dynamic contagion process
2.2.1 Time-homogeneous compound dynamic contagion process
For the standard DCP with time-homogeneous parameters, the infinitesimal generator $\cal A$} of the joint process $\{\left(\lambda_{t},N_{t},C_{t},M_{t},\Lambda_{t},t\right)\}_{t\in{\cal T}}$ acting on a function $f(\lambda,n,c,m,\Lambda,t)\in\mathcal{D}(\mathcal{A})$ is given by:
$\hskip-14.22636pt\mathcal{A}f(\lambda,n,c,m,\Lambda,t)$
@@ 2.2.1 Time-homogeneous compound dynamic contagion process
$\quad+\lambda\left[\int_{0}^{\infty}\int_{0}^{\infty}f(\lambda+y,n+1,c+\xi,m,\Lambda,t)\mathop{}\!\mathrm{d}G(y)\mathop{}\!\mathrm{d}J(\xi)-f(\lambda,n,c,m,\Lambda,t)\right]$
$\quad+\rho\left[\int_{0}^{\infty}f(\lambda+x,n,c,m+1,\Lambda,t)\mathop{}\!\mathrm{d}H(x)-f(\lambda,n,c,m,\Lambda,t)\right],$ (2.2)
where $\mathcal{D}(\mathcal{A})$ is the domain of the generator $\mathcal{A}$ such that $f(\lambda,n,c,m,\Lambda,t)$ is differentiable with respect to $\lambda$ and $t$, and
$\left|\int_{0}^{\infty}\int_{0}^{\infty}f(\lambda+y,n+1,c+\xi,m,\Lambda,t)\mathop{}\!\mathrm{d}G(y)\mathop{}\!\mathrm{d}J(\xi)-f(\lambda,n,c,m,\Lambda,t)\right|<\infty\text{,}$
$\left|\int_{0}^{\infty}f(\lambda+x,n,c,m+1,\Lambda,t)\mathop{}\!\mathrm{d}H(x)-f(\lambda,n,c,m,\Lambda,t)\right|<\infty\text{.}$
For details on the definition and derivations of the infinitesimal generator in (2.2.1), please refer to Dassios and Zhao (2011, 2017), Jang and Oh (2021) and Øksendal (2013).
Based on Section 2.2.1, we can easily obtain the moments of $\lambda_{t}$, $N_{t}$ and $C_{t}$. To do so, we denote the first-order moments of $X$, $Y$ and $\Xi$, respectively, by
@@
$\mathbb{E}\left[\lambda_{t}\mid\lambda_{0}\right]~=~\left\{\begin{array}[]{ll}\frac{\rho\mu_{H}+a\delta}{\kappa}+\left(\lambda_{0}-\frac{\rho\mu_{H}+a\delta}{\kappa}\right)\mathrm{e}^{-\kappa t},&\kappa\neq 0,\\ \lambda_{0}+\left(\rho\mu_{H}+a\delta\right)t,&\kappa=0,\end{array}\right$.
where $\kappa:=\delta-\mu_{G}$. Under the stationary condition $\kappa>0$, the asymptotic first moment of $\lambda_{t}$ under $\mathbb{P}$ is given by
$\mu_{1}\ :=\ \lim_{t\rightarrow\infty}\mathbb{E}\left[\lambda_{t}\mid\lambda_{0}\right]\ =\ \frac{\rho\mu_{H}+a\delta}{\kappa}.$
@@
$\mathbb{E}\left[N_{t}\mid\lambda_{0}\right]~=~\left\{\begin{array}[]{ll}\mu_{1}t+\left(\lambda_{0}-\mu_{1}\right)\frac{1}{\kappa}\left(1-\mathrm{e}^{-\kappa t}\right),&\kappa\neq 0,\\ \lambda_{0}t+\frac{1}{2}\left(\rho\mu_{H}+a\delta\right)t^{2},&\kappa=0.\end{array}\right$.
@@
$\mathbb{E}\left[C_{t}\mid\lambda_{0}\right]~=~\left\{\begin{array}[]{ll}\mu_{J}[\mu_{1}t+\left(\lambda_{0}-\mu_{1}\right)\frac{1}{\kappa}\left(1-\mathrm{e}^{-\kappa t}\right)],&\kappa\neq 0,\\ \mu_{J}[\lambda_{0}t+\frac{1}{2}\left(\rho\mu_{H}+a\delta\right)t^{2}],&\kappa=0.\end{array}\right$.
2.2.2 Time-inhomogeneous compound dynamic contagion process
The previously defined DCP $\lambda_{t}$ from Eq. 2.1 is the result of solving the stochastic differential equation (SDE)
$\lambda_{t}=\lambda_{0}+\int_{0}^{t}\delta\left(a-\lambda_{s}\right)\mathop{}\!\mathrm{d}s+\sum_{i\geq 1}X_{i}\,\mathbb{I}_{\{T_{1,i}\leq t\}}+\sum_{j\geq 1}Y_{j}\,\mathbb{I}_{\{T_{2,j}\leq t\}}\,.$ (2.4)
The parameters of this process are all time-homogeneous. More generally, we can extend to the time-inhomogeneous situation by allowing the parameters to vary over time (except the decaying rate $\delta$). Specifically, we impose the following assumptions:
  1. (A1)
    The mean-reverting level $a(t)$ and the rate of externally excited jump arrival $\rho(t)$ are deterministic functions of time $t$; they are bounded on all intervals $[0,t)$ (no explosions); and $\lambda_{0}\geq a(t)$.
  2. (A2)
    The distribution function of the externally excited jump sizes at any time $t$ is $H(x;t)$ for $x>0$ with $\mu_{H}(t)=\int_{0}^{\infty}x\,\mathop{}\!\mathrm{d}H(x;t)$.
  3. (A3)
    The distribution function of the self-excited jump sizes at any time $t$ is $G(y;t)$ for $y>0$ with $\mu_{G}(t)=\int_{0}^{\infty}y\,\mathop{}\!\mathrm{d}G(y;t)$.
  4. (A4)
    $a(t)$, $\rho(t)$, $G(y;t)$ and $H(x;t)$ are Riemann-integrable functions of $t$; they are all positive.
After replacing the time-homogeneous constants in Eq. 2.4 with time-varying functions, and solving the SDE, we arrive at the $\lambda_{t}$ for this new time-inhomogeneous DCP
$\lambda_{t}=\lambda_{0}\mathrm{e}^{-\delta t}+\delta\int_{0}^{t}a(s)\mathrm{e}^{-\delta(t-s)}\,\mathop{}\!\mathrm{d}s+\sum_{i\geq 1}X_{i}\mathrm{e}^{-\delta\left(t-T_{1,i}\right)}\mathbb{I}_{\{T_{1,i}\leq t\}}+\sum_{j\geq 1}Y_{j}\mathrm{e}^{-\delta\left(t-T_{2,j}\right)}\mathbb{I}_{\{T_{2,j}\leq t\}}\,.$ (2.5)
We allow the model parameters to be time-inhomogeneous or time-varying to reflect evolving exposure and claim dynamics in the insurance and reinsurance portfolios. Specifically, in reality, claim intensity and contagion effects may vary over time due to changes in insured exposure, portfolio composition, seasonal effects, and external risk conditions. Introducing time-inhomogeneity at this stage provides a flexible actuarial framework for modeling claim arrivals, while its role under an equivalent change of measures will become apparent later in Section 3.